OpenStudy (anonymous):

Medal Reward for help! Allie is designing a poster for her wall. Her design is a rectangle that is 8 inches wide and 12 inches long, but she wants to increase the size by a scale factor of 3 for display. What will the length of the poster be?

OpenStudy (anonymous):

It would be 2

OpenStudy (anonymous):

To increase a shape by a scale factor simply means to multiply the quantities. Currently the poster is 8x12. If you want to increase it by a factor of 3, then we just multiply 8*3 = 24, and 12*3 = 36. So the new dimensions should be 24x36. We can check this by dividing 12/8 = 3/2. And dividing 36/24 = 18/12 = 9/6 = 3/2. The proportions are the same!

OpenStudy (anonymous):

So to answer the question, since 12 was the length of the original poster, 36 should be the new length.

OpenStudy (anonymous):

@TurtleMuffin: Do you know the answer to this question: sin(x)=1/7

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

and don't call me dumb

OpenStudy (anonymous):

Unfortunately that can only be calculated with a calculator since arcsin(1/7) is 0.14334756... However, if the question asked to find cos(x) (or some other trig function), given that you know sin(x) = 1/7, then it can be done by drawing a picture. Actually, mostly anything that involves a trig function can be solved by drawing a triangle. You know that the sine of an angle is just the opposite over the hypotenuse. That's the definition of sine! So sin(x) = 1/7 simply says that for some triangle that has an angle 'x', the opposite side of 'x' is 1, and the hypotenuse of the triangle is 7. So now you can find the adjacent side of 'x', using the Pythagorean theorem.

OpenStudy (anonymous):

arcsin(x) is the inverse function of sin(x). It just undoes the operation. For example, arcsin(sin(x)) = x and sin(arcsin(x)) = x.

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (anonymous):

Can you just tell me what it is..lol..I don got time to work it out anyways..lol

OpenStudy (anonymous):

The answer to what? sin(x)=1/7, or the scale factor poster problem? The explanation to that is above. The new length is 36. Is the explanation not clear?

OpenStudy (anonymous):

oh ok thanks turtlemuffin

OpenStudy (anonymous):

You simply can't solve for x in sin(x) = 1/7 by hand lol. There has to be more to the problem.

OpenStudy (anonymous):

how do i give medels?

OpenStudy (anonymous):

*medals

OpenStudy (anonymous):

Is it 3.2, 8.2, 12.4, or 14.3 degrees.?? @TurtleMuffin

OpenStudy (anonymous):

U click the best response button@ OtakuGal

OpenStudy (anonymous):

Oh it would be 8.2 degrees. The original answer I gave is in radians. I guess they just wanted you to be able to use the arcsin(x) function. It should be a button somewhere on the calculator.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

What about this one @TurtleMuffin: On her way home, Carrie must drive up a hill that has an incline of 5 degrees If Carrie’s vehicle stops at an elevation 0.25 miles higher than where she started, how far did she drive up the hill? (This is a 30-60-90 Triangle by the way!) A. 0.25 miles B. 0.35 miles C. 2.87 miles D. 314.76 miles

OpenStudy (anonymous):

@TurtleMuffin where did u go?? I NEED your brains!!LOL

OpenStudy (anonymous):

i got some.

OpenStudy (anonymous):

Sorry, eh. I don't think it's a 30-60-90 triangle since the first sentence tells you the angle: it's 5 degrees. Pictures are always important:  I hope you can see how this picture is related to the problem. We're trying to solve for the distance of the hypotenuse, 'd'. We need to relate our angle, 5 degrees, and its opposite side and its hypotenuse. What trig function does that? The sine function! So, set it up: sin(5 degrees) = 0.25 / d. Solve for 'd' using algebra: d * sin(5 degrees) = 0.25 d = 0.25 / sin(5 degrees). Plugging that into a calculator one finds that the result is 2.868, approximately 2.87, so choice C.